# MRAG-Suite: A Cross-Domain Diagnostic Benchmark and Tooling for Visual Retrieval-Augmented Generation

### v2

@@ 1 Introduction
Existing multimodal QA datasets such as ScienceQA (Lu2022LearnTE), MMMU (Yue2023MMMUAM), and MMMU-Pro (Yue2024MMMUProAM) supply the necessary visual context in the input, side-stepping retrieval. Retrieval-centric efforts like Visual Haystacks (Wu2024VisualHA) emphasize hard look-up but focus on a single domain and lack fine-grained diagnosis. Text-focused diagnostic tools (e.g., RAGChecker, RefChecker) (ru2024ragchecker; Hu2024RefCheckerRF) provide claim-level analysis, but cannot assess how claims are grounded in visual evidence.
@@
    Claim-level multimodal diagnostics. MM-RAGChecker extends textual claim checking (ru2024ragchecker; Hu2024RefCheckerRF) to visual contexts, attributing support or errors to specific modalities and retrieved items.
@@ 2 Related Work
Traditional evaluation of RAG systems primarily relied on metrics like exact match and ROUGE scores, which are insufficiently fine-grained to diagnose specific failures. RAGChecker (ru2024ragchecker) introduced claim-level diagnostics for textual RAG, evaluating retrieval accuracy, factual consistency, and generation hallucinations. However, these tools remain limited to text-only contexts, lacking the capability to assess visual evidence integration and multimodal reasoning.
@@
Benchmarks such as ScienceQA (Lu2022LearnTE), MMMU (Yue2023MMMUAM), and their advanced variants MMMU-Pro (Yue2024MMMUProAM) evaluate models’ abilities to reason jointly over provided image and text inputs but exclude retrieval challenges. In contrast, Visual Haystacks (Wu2024VisualHA) targets multimodal retrieval tasks but is constrained to single-domain visual reasoning, omitting extensive claim-level diagnostics and failing to capture multimodal reasoning intricacies across diverse domainsli2025human; li2025chatmotion.
@@
Recent studies reveal significant biases in multimodal RAG systems. For example, Yao et al. (Yao2025WhoII) demonstrated strong positional and modality biases, with models disproportionately relying on initial evidence items or textual over visual modalities. However, existing benchmarks have not systematically quantified these biases across multiple domains under controlled experimental conditions. MRAG-Suite specifically addresses this shortcoming by incorporating varied distractor and ambiguity scenarios to measure these biases explicitlywang2025selfdestructivelanguagemodel; liang2025graphrag; jiang2024robustkv; dan2024evaluation; lu2025predicting; li2025frequency.
@@
Prior work on multimodal QA has predominantly focused on specific domains or modalities, such as ChartMRAG (yang2025benchmarking) and VDocRAG (Tanaka2025VDocRAGRG), demonstrating the feasibility and utility of retrieval-augmented QA systems in targeted contexts. Nonetheless, these domain-specific evaluations lack a unified framework for cross-domain comparisons and fine-grained diagnostics across modalities, a gap that MRAG-Suite explicitly aims to fill.
@@ 3.2 Corpora Overview
  • WebQA chang2022WebQA: open-web questions with both images (captions/descriptions) and short text snippets; often multi-hop and open-ended.
  • Chart-RAG yang2025benchmarking(from ChartQA masry2022chartqa): chart understanding (e.g., “When did X peak?”); evidence includes the chart image plus OCR/metadata; answers are numeric/categorical values.
  • Visual-RAG Wu2025VisualRAGBT: queries that require visual knowledge (object/attribute recognition), drawn from Visual7W/FVQA/MRAG-Bench-style scenarios.
  • MRAG-Bench hu2024mrag: 1,353 vision-centric MCQ converted to open-ended by taking the correct option as $a^{\text{short}}$; emphasizes fine-grained and multi-view recognition.
  • VisRAG Yu2024VisRAGVR subsets: Arxiv Li2024MultimodalAA (scholarly PDF pages), Plot Methani2019PlotQARO (technical plots), Slide 10.1609/aaai.v37i11.26598 (slide decks), and Document Tito2022HierarchicalMT (scanned manuals/forms). Each page is treated as a visual document with OCR text.
@@ 3.3 Normalization and Indexing
All samples are normalized to {question, short_answer, long_answer, evidence_imgs, evidence_txts}. Images use CLIP (ViT-L/14) features plus captionsradford2021learning; text passages use dense embeddings. Visual items are indexed in FAISS; text is stored in a DPR/ElasticSearch store.

### v3

@@ 1 Introduction
Existing multimodal QA datasets such as ScienceQA (Lu et al., 2022), MMMU (Yue et al., 2023), and MMMU-Pro (Yue et al., 2024) supply the necessary visual context in the input, side-stepping retrieval. Retrieval-centric efforts like Visual Haystacks (Wu et al., 2024) emphasize hard look-up but focus on a single domain and lack fine-grained diagnosis. Text-focused diagnostic tools (e.g., RAGChecker, RefChecker) (Ru et al., 2024; Hu et al., 2024b) provide claim-level analysis, but cannot assess how claims are grounded in visual evidence.
@@
    Claim-level multimodal diagnostics. $M$-RAGChecker extends textual claim checking (Ru et al., 2024; Hu et al., 2024b) to visual contexts, attributing support or errors to specific modalities and retrieved items.
@@ 2 Related Work
Traditional evaluation of RAG systems primarily relied on metrics like exact match and ROUGE scores, which are insufficiently fine-grained to diagnose specific failures. RAGChecker (Ru et al., 2024) introduced claim-level diagnostics for textual RAG, evaluating retrieval accuracy, factual consistency, and generation hallucinations. However, these tools remain limited to text-only contexts, lacking the capability to assess visual evidence integration and multimodal reasoning.
@@
Benchmarks such as ScienceQA (Lu et al., 2022), MMMU (Yue et al., 2023), and their advanced variants MMMU-Pro (Yue et al., 2024) evaluate models’ abilities to reason jointly over provided image and text inputs but exclude retrieval challenges. In contrast, Visual Haystacks (Wu et al., 2024) targets multimodal retrieval tasks but is constrained to single-domain visual reasoning, omitting extensive claim-level diagnostics and failing to capture multimodal reasoning intricacies across diverse domainsLi et al. (2025b, a).
@@
Recent studies reveal significant biases in multimodal RAG systems. For example, Yao et al. (Yao et al., 2025) demonstrated strong positional and modality biases, with models disproportionately relying on initial evidence items or textual over visual modalities. However, existing benchmarks have not systematically quantified these biases across multiple domains under controlled experimental conditions. MRAG-Suite specifically addresses this shortcoming by incorporating varied distractor and ambiguity scenarios to measure these biases explicitlyWang et al. (2025); Liang et al. (2025); Jiang et al. (2024); Dan et al. (2024); Lu et al. (2025); Li et al. .
@@
Prior work on multimodal QA has predominantly focused on specific domains or modalities, such as ChartMRAG (Yang et al., 2025) and VDocRAG (Tanaka et al., 2025), demonstrating the feasibility and utility of retrieval-augmented QA systems in targeted contexts. Nonetheless, these domain-specific evaluations lack a unified framework for cross-domain comparisons and fine-grained diagnostics across modalities, a gap that MRAG-Suite explicitly aims to fill.
@@ 3.2 Corpora Overview
  • WebQA Chang et al. (2022): open-web questions with both images (captions/descriptions) and short text snippets; often multi-hop and open-ended.
  • Chart-RAG Yang et al. (2025)(from ChartQA Masry et al. (2022)): chart understanding (e.g., “When did X peak?”); evidence includes the chart image plus OCR/metadata; answers are numeric/categorical values.
  • Visual-RAG Wu et al. (2025): queries that require visual knowledge (object/attribute recognition), drawn from Visual7W/FVQA/MRAG-Bench-style scenarios.
  • MRAG-Bench Hu et al. (2024a): 1,353 vision-centric MCQ converted to open-ended by taking the correct option as $a^{\text{short}}$; emphasizes fine-grained and multi-view recognition.
  • VisRAG Yu et al. (2024) subsets: Arxiv Li et al. (2024) (scholarly PDF pages), Plot Methani et al. (2019) (technical plots), Slide Tanaka et al. (2023) (slide decks), and Document Tito et al. (2022) (scanned manuals/forms). Each page is treated as a visual document with OCR text.
@@ 3.3 Normalization and Indexing
All samples are normalized to {question, short_answer, long_answer, evidence_imgs, evidence_txts}. Images use CLIP (ViT-L/14) features plus captionsRadford et al. (2021); text passages use dense embeddings. Visual items are indexed in FAISS; text is stored in a DPR/ElasticSearch store.